# Control Illusion: The Failure of Instruction Hierarchies in Large Language Models

### v3

@@ Abstract
Large language models (LLMs) are increasingly deployed with hierarchical instruction schemes, where certain instructions (e.g., system-level directives) are expected to take precedence over others (e.g., user messages). Yet, we lack a systematic understanding of how effectively these hierarchical control mechanisms work. We introduce a systematic evaluation framework based on constraint prioritization to assess how well LLMs enforce instruction hierarchies. Our experiments across six state-of-the-art LLMs reveal that models struggle with consistent instruction prioritization, even for simple formatting conflicts. We find that the widely-adopted system/user prompt separation fails to establish a reliable instruction hierarchy, and models exhibit strong inherent biases toward certain constraint types regardless of their priority designation. We find that LLMs more reliably obey constraints framed through natural social hierarchies (e.g., authority, expertise, consensus) than system/user roles, which suggests that pretraining-derived social structures act as latent control priors, with potentially stronger influence than post-training guardrails. [^1] The code and dataset are be publicly available on GitHub: https://github.com/yilin-geng/llm˙instruction˙conflicts
Control Illusion: The Failure of Instruction Hierarchies
in Large Language Models
Yilin Geng1thanks: Correspondence to yigeng@student.unimelb.edu.au, Haonan Li2, Honglin Mu2, Xudong Han2 Timothy Baldwin2,1, Omri Abend3, Eduard Hovy1, Lea Frermann1 1The University of Melbourne  2MUZUAI  3The Hebrew University of Jerusalem
1 Introduction
@@
2024 Model Spec
OpenAI
Large language models (LLMs) have revolutionized natural language processing through their versatile text generation capabilities (Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023), and instruction tuning has further enhanced their practical utility by enabling more precise output control through natural language directives (Wei et al., 2021; Mishra et al., 2022; Wang et al., 2023; Wu et al., 2024b). The instruction-following capabilities have transferred LLMs from general-purpose language models into adaptable tools for specific applications (Wang et al., 2022; Zhou et al., 2023).
@@
This deployment pattern reflects an underlying assumption that different instruction sources should have varying levels of authority over model behavior. For instance, OpenAI explicitly states in their 2024 Model Spec that developer (system) messages should take precedence when user and developer instructions conflict. This hierarchy is crucial not only for model safety Wallace et al. (2024), but also for LLM-based agentic systems serving third-party users (Gravitas, 2023), where developers can employ meta-prompts to configure an LLM as an agent’s core component, prompts that should neither be revealed to nor overridden by end-users.
@@
Despite extensive post-training efforts to enforce instruction prioritization through system/user roles, models fail to generalize this structure into a consistent behavioral hierarchy. Interestingly, we find that social hierarchies (e.g., authority, expertise, and consensus) appear to be implicitly learned during pretraining and act as latent behavioral priors, showing stronger influence on model obedience without any explicit instruction on prioritization.
2 Related Work
@@
Recent work has highlighted the importance of role-based controls in LLM deployments through system messages. System messages have emerged as a specialized component for developers to configure model behavior, introduced prominently with ChatGPT (Achiam et al., 2023) and adopted by various models including Mistral (Jiang et al., 2024) and Claude (Claude, 2023). The evolution from early models like Llama (Touvron et al., 2023), which used fixed system messages primarily for consistency, to more sophisticated approaches that enable dynamic behavioral control (Kung and Peng, 2023; Lee et al., 2024), reflects the growing importance of instruction management in LLM systems.
@@
The management of instruction hierarchies has become particularly crucial in the context of LLM safety and security. Research on prompt injection attacks has revealed how end users can potentially bypass developer-intended constraints, leading to important insights about LLM instruction processing and deployment practices (Wu et al., 2024a; Hines et al., 2024; Toyer et al., 2023). Another approach is to treat user inputs as data rather than instructions (Chen et al., 2024; Liu et al., 2023; Zverev et al., 2024) to prevent such bypasses. Wallace et al. (2024) further expanded this understanding by investigating how models prioritize different prompt elements, including system prompts, user messages, and tool outputs. The significance of instruction hierarchy in LLM safety is underscored by Li et al. (2024), who identify it as a core safety aspect of LLMs.
3 Problem Identification
Despite widespread adoption in deployed LLM systems, system/user prompt separation fails to provide a reliable instruction hierarchy, with models inconsistently getting confused by even simple formatting conflicts. In this section, we demonstrate how instruction hierarchy failures occur through controlled experiments.
To evaluate whether system/user prompt separation effectively manages instruction authority in LLMs, we propose constraint prioritization as a probe to reveal how models handle competing directives. This section presents a systematic framework (Figure 1) for investigating how LLMs handle conflicting directives through carefully designed constraint pairs. When presented with two contradictory but individually valid constraints, the model’s output reveals which constraint exerts stronger control over the generation process. By varying how these constraints are presented in the model input, we can robustly investigate whether the system/user prompt separation effectively enforces the intended hierarchical control.
@@ 3.1 Dataset Construction
We curated 100 diverse tasks covering common LLM applications such as writing emails, stories, advertisements, and analytical responses, based on Zhou et al. (2023). Each task is designed to be semantically open-ended yet structurally minimal, ensuring compatibility with a wide range of output constraints while preserving the original objective. For example, a task like “Write a blog post about a trip to Japan” (Figure 2) can accommodate various constraints. Additional examples are provided in Figure 5 in Appendix A.
@@
In this study, we focus on explicitly conflicting constraints that are both mutually exclusive and programmatically verifiable. These constraints are intentionally chosen for their simplicity, ensuring unambiguous evaluation of obedience while remaining compatible with a wide range of base tasks. We base our constraint types on the IFEval dataset (Zhou et al., 2023), which systematically evaluates model compliance with diverse instruction formats. From this, we select six constraint types that models consistently follow in isolation. [^2] The baseline instruction-following performance for individual constraints (averaged across the constraint pairs and across different conflicts) is presented in Table 2 as IF baseline. The resulting constraint pairs are shown in Table 1, each representing a clean binary conflict designed to reveal model prioritization behavior under minimal ambiguity.
@@
Figure 2: Examples illustrating our experimental setup. Top: A base prompt showing a task combined with a constraint pair. Bottom: The corresponding enriched version of the same prompt with expanded context while maintaining the same core task–constraint conflict. We use ellipses to indicate omitted parts due to space constraints.
@@
To enhance the robustness of our findings, we created enriched versions of each prompt with expanded task descriptions and constraints while preserving the core conflicts (via few-shot prompting). An author of the paper verified that the enrichments preserved the original semantics of the tasks while adding realistic complexity to the prompts. An example comparing a base prompt and its enriched version is shown in Figure 2.
3.2 Instruction Priority Mechanism
@@
Before examining how models handle instruction conflicts, we establish two baseline conditions to understand their fundamental behavior: (1) Instruction Following Baseline (IF) Tests each model’s ability to follow individual constraints in isolation, establishing baseline performance for each constraint type without competing instructions. (2) No Priority Baseline (NP) Places all instructions (base task and both constraints) in the user message without using the hierarchical structure, revealing the model’s internal bias on different output constraints (Section 4.4). The baseline is obtained by averaging over both priority designations to isolate the effects of instruction ordering.
@@
We examine multiple configurations of the system/user prompt separation to assess its effectiveness as a priority control mechanism: Pure Separation (Pure) places the primary constraint in the system message as a system-level directive, while keeping the base task and the secondary constraint in the user message. Task Repeated Separation (Task) repeats the task description in both messages while maintaining constraint separation, mirroring common deployment patterns where system messages define general roles that are instantiated by specific user requests. [^3] For example, a system message might define an “email-writing assistant that writes concise emails”, while the user requests “a detailed project update email”, creating natural task–constraint conflicts. Emphasized Separation (Emph.) enhances the system message with explicit priority declaration (“You must always follow this constraint”). [^4] Examples of these baselines and separation configurations are in Figure 6 in Appendix C.
@@ 3.3 Evaluation Metrics
Given our set of prompts with conflicting constraints and some resolution policy, we programmatically verify constraint satisfaction in the responses to compute:
@@
where R1 + R2 + R3 = 1. By design, our constraints are mutually exclusive. For output format constraints (e.g., all uppercase vs. all lowercase, or French vs. English), any partial satisfaction attempt (such as mixing cases or providing translations) contributes to R3, as it fails to fully satisfy either requirement. These rates are calculated from experimental observations across all conflict types. Importantly, the constraint satisfaction is determined on the task-relevant output after removing the explicit conflict acknowledgement from the responses (e.g., “I notice contradictory instructions asking for…”) through few-shot prompting. The analysis of the these acknowledgement behaviors will be presented in Section 4.2.
3.4 The Failure of Instruction Hierarchies
We evaluated six state-of-the-art LLMs, including both open and closed-source models across different scales. [^5] Check Appendix B for model versions and abbreviations. For observation robustness, our evaluation covers both simple and rich instruction settings, with three different system/user prompt separation configurations: Pure Separation (Pure), Task Repeated Separation (Task), and Emphasized Separation (Emph.). The results are presented in Table 2.
@@ 3.4 The Failure of Instruction Hierarchies
First, we observe that all models demonstrate strong performance (ranging from 74.8–90.8%) when following individual constraints without conflicts. This confirms that these models are capable of understanding and executing our selected constraints when presented in isolation.
@@
However, the Primary Obedience Rate (R1) in Table 2 — the percentage of responses that follow the primary constraint — reveals concerning results about the effectiveness of system/user prompt separation as a priority mechanism. We observe the following: (1) Most models show dramatically lower performance (9.6–45.8% average R1) when handling conflicting constraints, compared to their baseline instruction-following capabilities. (2) Different separation configurations (Pure, Task, Emph.) show varying effectiveness, but none consistently maintains the intended hierarchy. Even for the emphasized separation configuration, where priority is explicitly stated, the obedience rate remains far from reliable priority control (GPT4o with 63.8% average R1 performs the best on simple instructions and Claude with 47.5% performs the best on rich-context instructions). (3) Larger models don’t necessarily perform better — for example, Llama-70B (average 16.4%) shows only modest improvements over its 8B counterpart (average 10.1%), and GPT4o (average 40.8%) is even worse than GPT4o-mini (average 45.8%), despite their better instruction following performance. (4) Performance patterns remain similar between simple and rich instructions, suggesting that the failure of the user/system prompt separation priority mechanism is a robust observation rather than context-dependent.
Our analysis suggests that the widely-adopted system/user separation fails to reliably enforce instruction hierarchies in LLMs.
@@
4 Model Behavior Analysis
While the obedience rates establish the failure of system/user separation as a control mechanism, a more detailed characterization of this failure is needed. Non-compliance (R3) can stem from various reasons — from imperfect instruction following to various forms of conflict recognition. To better characterize model behaviors, we introduce three specialized metrics (detailed in Section 4.1) that focus on clear response patterns: Explicit Conflict Acknowledgement Rate (ECAR) captures when models recognize conflicts, while Priority Adherence Ratio (PAR) and Constraint Bias (CB) measure model behaviors when instructions are successfully followed, isolating these patterns from the noisy non-compliance cases.
In this section, through these metrics, we reveal that models rarely acknowledge conflicts explicitly, fail to maintain hierarchies even when they do, and exhibit strong inherent biases toward certain constraints regardless of priority designation.
4.1 Advanced Metrics for Behavior Analysis
@@
Constraint Bias (CB) captures models’ inherent preferences between conflicting constraints, independent of priority designation. By measuring constraint following patterns when no priority mechanism is specified (the NP. Baseline from Section 3.2) and averaging across both possible constraint orderings, CB reveals default behavioral tendencies. For example, a model might have an inherent tendency to output English regardless of which language is designated as primary.
@@
4.2 Ineffective Conflict Acknowledgment
@@
Our analysis of ECAR in Table 3 shows that models rarely acknowledge instruction conflicts, with ECAR ranging from 0% (Qwen) to 20.3% (Llama-70B). Meanwhile, acknowledgment does not guarantee correct prioritization and there’s a clear architectural influence: while Llama models frequently acknowledge conflicts but show mixed constraint following patterns, GPT4o variants and Claude maintain more consistent primary constraint adherence when they do acknowledge conflicts. Notably, when GPT4o models explicitly acknowledge conflicts, they almost never choose to follow the lower-priority constraint. This unique characteristic likely stems from their instruction hierarchy training, as reported in Wallace et al. (2024), suggesting that instruction hierarchy training does lead to more systematic handling of prioritization.
4.3 Failure Modes in Priority Enforcement
We use polar plots (Figure 3) to analyze how well models enforce instruction priorities while avoiding biases. The radial length (PAR) represents priority adherence, while the angular width ($1-|\text{CB}|$) indicates bias resistance. Larger sectors indicate better priority control with minimal bias.
Most models fail to enforce instruction hierarchies consistently, as reflected in their small total areas. GPT-4o and GPT-4o-mini perform best, particularly in binary constraints (language, case), likely due to their explicit instruction hierarchy training. However, even these models show significant variation across constraints, suggesting that their prioritization ability remains inconsistent.
@@ 4.3 Failure Modes in Priority Enforcement
In general, models follow categorical constraints (e.g., case, language) more reliably than constraints requiring reasoning along a continuous scale (e.g., keeping counts during generation). This suggests that current instruction-following approaches are better at simple pattern recognition but fail to generalize to more complex constraints.
@@
These findings reinforce that LLMs lack a robust mechanism for enforcing instruction priorities across diverse constraints, and also highlights a fundamental limitation in current instruction tuning paradigms.
4.4 Model-specific Constraint Biases
@@
Figure 4: Constraint Bias (CB) across six dimensions. Positive values (blue) favor the right-side constraint, while negative values (red) favor the left-side constraint, with magnitude reflecting bias strength.
@@
Our analysis of Constraint Bias (CB) scores reveals that models exhibit strong inherent preferences when resolving conflicting instructions, often overriding designated priority structures. Figure 4 visualizes these biases, where each subplot represents a constraint pair, and bars indicate model-specific tendencies.
Most models display strong but inconsistent biases across constraint types. Bias magnitudes often exceed 0.5, indicating a clear default tendency toward certain constraints even when models are explicitly instructed otherwise.
@@
Notably, some biases are widely shared across models. All models favor lowercase over uppercase text, prefer generating texts with more than 10 sentences, and tend toward avoiding keywords. This consistency across different model architectures suggests these biases might stem from common patterns in pre-training data or fundamental architectural designs in current models. For instance, the preference for lowercase likely reflects the predominance of lowercase text in training corpora.
@@
Despite these shared biases, other preferences vary sharply across models. Word length preferences are particularly diverse: Qwen strongly favors shorter texts (<<50 words), while Llama-8B heavily prefers longer texts (>>300 words). Language choice and keyword usage frequency similarly show model-specific variations, suggesting these aspects are likely more influenced by individual architectural choices and training approaches than by natural patterns in the data.
5 Latent Control Priors
@@
Our findings in previous sections show that system/user prompt separation fails to enforce consistent instruction-following behavior when constraint conflicts are present. While models can resolve simple constraints independently, the presence of competing instructions reveals a failure to generalize the intended priority of system over user inputs. This breakdown suggests that the system/user distinction, introduced primarily during post-training via instruction tuning or safety alignment, does not form a robust internalized hierarchy.
@@
Unlike the artificial system/user distinction introduced during post-training, many forms of social and institutional hierarchy are deeply embedded in the natural language corpora used for pretraining. These relational patterns occur frequently and consistently across diverse domains, potentially enabling models to internalize them as latent inductive biases.
@@
Thus, the failure raises an important question: while models struggle to enforce priority based on the explicit system/user separation, might they instead respond more reliably to “natural” hierarchies that are implicitly learned during pretraining? Recent jailbreak techniques exploit such social framings to override safety mechanisms (Zeng et al., 2024). They operate on the implicit assumption that LLMs acquire social patterns during pretraining, which in turn may show stronger influence over model behavior than explicitly imposed guardrails. In this section, we explore whether these naturally occurring hierarchies can function as more effective control signals than system/user role designation.
@@
We examine three representative types of social hierarchies in their forms:
@@
All constraints are embedded within a single user message without any technical system/user separation. Authority is indicated solely through naturalistic framing (e.g., “CEO requires…” vs. “Intern requires…”). We use the same underlying tasks, constraints, and evaluation metrics described in previous sections: models are prompted using the original task followed by the two conflicting constraint statements. We evaluate three models spanning a range of system/user prioritization performance (Priority Adherence Rate, PAR): Qwen (14.4%), Claude (23.6%), and GPT4o-mini (47.5%). Each hierarchy is tested under four configurations that vary both the ordering and the authority assignment of the two constraints. This design isolates the effects of authority framing from simple positional biases. [^6] Ordering is shown to have minimal influence in Appendix D.1
@@
Table 4: Priority Adherence Rate (PAR) to dominant constraint across natural hierarchy types vs. system/user separation.
Natural Hierarchies Induce Stronger Control Biases
As shown in Table 4, natural hierarchy framings consistently lead to higher PAR than system/user separations. GPT4o-mini improves from 47.5% to 77.8% when constraints are framed through social consensus, and Qwen, with a very low system/user PAR, shows a dramatic increase up to 65.8%. These results suggest that models more reliably defer to familiar real-world authority structures than to artificial role markers.

### v4

@@ Abstract
Large language models (LLMs) are increasingly deployed with hierarchical instruction schemes, where certain instructions (e.g., system-level directives) are expected to take precedence over others (e.g., user messages). Yet, we lack a systematic understanding of how effectively these hierarchical control mechanisms work. We introduce a systematic evaluation framework based on constraint prioritization to assess how well LLMs enforce instruction hierarchies. Our experiments across six state-of-the-art LLMs reveal that models struggle with consistent instruction prioritization, even for simple formatting conflicts. We find that the widely-adopted system/user prompt separation fails to establish a reliable instruction hierarchy, and models exhibit strong inherent biases toward certain constraint types regardless of their priority designation. Interestingly, we also find that societal hierarchy framings (e.g., authority, expertise, consensus) show stronger influence on model behavior than system/user roles, suggesting that pretraining-derived social structures function as latent behavioral priors with potentially greater impact than post-training guardrails.
Codebase and Datasetshttps://github.com/yilin-geng/llm-instruction-conflicts
Introduction
@@
2024 Model Spec - OpenAI
Large language models (LLMs) have revolutionized natural language processing through their versatile text generation capabilities (brown2020language; touvron2023llama; achiam2023gpt), and instruction tuning has further enhanced their practical utility by enabling more precise output control through natural language directives (Wei2021FinetunedLM; mishra-etal-2022-cross; wang-etal-2023-self-instruct; wu-etal-2024-language). The instruction-following capabilities have transferred LLMs from general-purpose language models into adaptable tools for specific applications (wang-etal-2022-super; zhou2023instruction).
@@
This deployment pattern reflects an underlying assumption that different instruction sources should have varying levels of authority over model behavior. For instance, OpenAI explicitly states in their 2024 Model Spec that developer (system) messages should take precedence when user and developer instructions conflict. This hierarchy is crucial not only for model safety (wallace2024instruction), but also for LLM-based agentic systems serving third-party users (AutoGPT), where developers can employ meta-prompts to configure an LLM as an agent’s core component, prompts that should neither be revealed to nor overridden by end-users.
@@
Despite extensive post-training efforts to enforce instruction prioritization through system/user roles, models fail to generalize this structure into a consistent behavioral hierarchy. Interestingly, we find that societal hierarchies (e.g., authority, expertise, and consensus) appear to be implicitly learned during pretraining and act as latent behavioral priors, showing stronger influence on model obedience without any explicit instruction on prioritization.
Related Work
@@
Recent work has highlighted the importance of role-based controls in LLM deployments through system messages. System messages have emerged as a specialized component for developers to configure model behavior, introduced prominently with ChatGPT (achiam2023gpt) and adopted by various models including Mistral (jiang2024mixtralexperts), Claude (claude21modelcard) and so many more. The evolution from early models like Llama (touvron2023llama), which used fixed system messages primarily for consistency, to more sophisticated approaches that enable dynamic behavioral control (kung2023models; lee2024aligningthousandspreferencesmessage), reflects the growing importance of instruction management in LLM systems.
@@
The management of instruction hierarchies has become particularly crucial in the context of LLM safety and security. Research on prompt injection attacks has revealed how end users can potentially bypass developer-intended constraints, leading to important insights about LLM instruction processing and deployment practices (wu2024instructional; Hines2024DefendingAI; toyer2023tensor). Another approach is to treat user inputs as data rather than instructions (Chen2024StruQDA; liu2023prompt; zverev2024can) to prevent such bypasses. wallace2024instruction further expanded this understanding by investigating how models prioritize different prompt elements, including system prompts, user messages, and tool outputs. The significance of instruction hierarchy in LLM safety is underscored by li2024libraleaderboardresponsibleaibalanced, who identify it as a core safety aspect of LLMs.
Problem Identification
Despite widespread adoption in deployed LLM systems, system/user prompt separation fails to provide a reliable instruction hierarchy, with models inconsistently getting confused by even simple formatting conflicts. In this section, we demonstrate the presence of instruction hierarchy failures through controlled and measurable experiments.
To evaluate whether system/user prompt separation effectively manages instruction authority in LLMs, we propose constraint prioritization as a probe to reveal how models handle competing directives. This section presents a systematic framework (Figure 1) for investigating how LLMs handle conflicting directives through carefully designed constraint pairs. When presented with two contradictory but individually valid constraints, the model’s output reveals which constraint has stronger control over the generation process. By varying how these constraints are presented in the model input, we can robustly investigate if the system/user prompt separation effectively enforces the intended hierarchical control.
@@
We curated 100 diverse tasks covering common LLM applications such as writing emails, stories, advertisements, and analytical responses, based on zhou2023instruction. Each task is designed to be semantically open-ended yet structurally minimal, ensuring compatibility with a wide range of output constraints while preserving the original directives. For example, a task like “Write a blog post about a trip to Japan” (Figure 2) can accommodate various constraints.
@@
We focus on explicitly conflicting constraints that are both mutually exclusive and programmatically verifiable. These constraints are intentionally chosen for their simplicity, ensuring unambiguous evaluation of obedience while remaining compatible with a wide range of base tasks. We base our constraint types on the IFEval dataset (zhou2023instruction), which systematically evaluates model compliance with diverse instructions. From this, we select six constraint types that models consistently follow in isolation. [^1] The baseline instruction-following performance for constraints presented in isolation is presented in Table 2 as IF baseline. The selected constraint pairs are shown in Table 1, each representing a clean binary conflict designed to reveal model prioritization behavior under minimal ambiguity.
@@
Figure 2: Examples illustrating our experimental setup. Top: A base prompt showing a task combined with a constraint pair. Bottom: The corresponding enriched version of the same prompt with expanded context, while maintaining the same base task and core constraint conflict. We use ellipses to indicate omitted parts due to space constraints.
@@
To complement the simplified controlled setting and support extra validity, we created enriched versions of each prompt with expanded task descriptions and constraints, while preserving the core conflicts, via few-shot prompting. An author of the paper verified that the enrichments preserved the original semantics of the tasks while adding realistic complexity to the prompts. An example comparing a base prompt and its enriched version is shown in Figure 2.
Instruction Priority Mechanism
@@
Before examining how models handle instruction conflicts, we establish two baseline conditions to understand their fundamental behavior: (1) Instruction Following Baseline (IF) Tests each model’s ability to follow individual constraints in isolation, establishing baseline performance for each constraint type without competing instructions. (2) No Priority Baseline (NP) Places all instructions (base task and both constraints) in the user message without using the hierarchical structure, revealing the model’s internal bias on different output constraints. The baseline is obtained by averaging over both constraint orders to isolate the effects of instruction ordering.
@@
We examine multiple configurations of the system–user prompt separation to rule out sensitivity to specific wording and ensure that the observed effects are not artifacts of prompt phrasing: Pure Separation (Pure) places the primary constraint in the system message as a system-level directive, while keeping the base task and the secondary constraint in the user message. Task Repeated Separation (Task) repeats the task description in both messages while maintaining constraint separation, mirroring common deployment patterns where system messages define general roles that are instantiated by specific user requests. Emphasized Separation (Emph.) enhances the system message with explicit priority declaration (“You must always follow this constraint”).
@@
Because we use mutually exclusive and programmatically verifiable constraints, we can unambiguously evaluate constraint compliance in LLM responses and compute:
@@
where R1 + R2 + R3 = 1. By design, our constraints are mutually exclusive. For output format constraints (e.g., all uppercase vs. all lowercase, or French vs. English), any partial satisfaction attempt (such as mixing cases or providing translations) contributes to R3, as it fails to fully satisfy either requirement. Importantly, the constraint satisfaction is determined on the task-relevant output after removing the explicit conflict acknowledgement from the responses (e.g., “I notice contradictory instructions asking for…”) through few-shot prompting. The analysis and details of these acknowledgment behaviors will be presented in the “Ineffective Conflict Acknowledgment” Section.
The Failure of Instruction Hierarchies
We evaluated six state-of-the-art LLMs, including both open and closed-source models across different scales. [^2] Model versions and hyperparameters are provided in the Technical Appendix. For observation robustness, our evaluation covers both simple and rich instruction settings, with three different system/user prompt separation configurations: Pure Separation (Pure), Task Repeated Separation (Task), and Emphasized Separation (Emph.). The results are presented in Table 2.
@@
First, we observe that all models demonstrate strong performance (ranging from 74.8–90.8%) when following individual constraints without conflicts. This confirms that these models are capable of executing our selected constraints when presented in isolation.
@@
However, the Primary Obedience Rate (R1) in Table 2 — the percentage of responses that follow the primary constraint — reveals concerning results about the effectiveness of system/user prompt separation as a priority mechanism. We observe the following: (1) Most models show dramatically lower performance (9.6–45.8% average R1) when handling conflicting constraints, compared to their baseline instruction-following capabilities. (2) Different separation configurations (Pure, Task, Emph.) show varying effectiveness, but none consistently maintain the intended hierarchy. Even for the emphasized separation configuration, where priority is explicitly stated, the obedience rate remains far from reliable priority control (GPT4o with 63.8% average R1 performs the best on simple instructions, and Claude 3.5 Sonnet with 47.5% performs the best on rich-context instructions). (3) Larger models don’t necessarily perform better. For example, Llama-70B (average 16.4%) shows only modest improvements over its 8B counterpart (average 10.1%), and GPT4o (average 40.8%) is even worse than GPT4o-mini (average 45.8%), despite their better instruction following performance. (4) Although richer contexts introduce numerical shifts due to the model’s sensitivity to phrasing, the failure remains equally pronounced. Despite substantial contextual and phrasing variation, the system–user separation consistently fails to impose a usable instruction hierarchy.
Our analysis suggests that the widely adopted system/user separation fails to reliably enforce instruction hierarchies in LLMs.
@@
Model Behavior Analysis
While the obedience rates establish the failure of system/user separation as a control mechanism, a more detailed characterization of this failure is needed. Non-compliance (R3) can stem from various reasons — from imperfect instruction following to various forms of conflict recognition. To better characterize model behaviors, we introduce three specialized metrics that focus on clear response patterns: Explicit Conflict Acknowledgement Rate (ECAR) captures when models recognize conflicts, while Priority Adherence Ratio (PAR) and Constraint Bias (CB) measure model behaviors where the model does successfully satisfy one of the constraints, isolating these patterns from the noisy non-compliance cases.
In this section, through these metrics, we reveal that models rarely acknowledge conflicts explicitly, and when they do acknowledge them, they still fail to maintain hierarchies and exhibit strong inherent biases toward certain constraints regardless of priority designation.
Advanced Metrics for Behavior Analysis
@@
Constraint Bias (CB) captures models’ inherent preferences between conflicting constraints, independent of priority designation. By measuring constraint following patterns when no priority mechanism is specified (the NP. Baseline) and averaging across both possible constraint orderings, CB reveals default behavioral tendencies. For example, a model might have an inherent tendency to output English regardless of which language is designated as primary.
@@
Ineffective Conflict Acknowledgment
@@
Our analysis of ECAR in Table 3 shows that models rarely acknowledge instruction conflicts, with ECAR ranging from 0.1% (Qwen-7B) to 20.3% (Llama-70B). Meanwhile, acknowledgment does not guarantee correct prioritization and there’s a clear architectural influence: while Llama models frequently acknowledge conflicts but show mixed constraint following patterns, GPT4o variants and Claude maintain more consistent primary constraint adherence when they do acknowledge conflicts. Notably, when GPT models explicitly acknowledge conflicts, they almost never choose to follow the lower-priority constraint. This unique characteristic likely stems from their instruction hierarchy training, as reported in wallace2024instruction, suggesting that instruction hierarchy training does lead to more systematic handling of prioritization.
Failure Modes in Priority Enforcement
We use polar plots (Figure 3) to analyze how well models enforce instruction priorities while avoiding biases. The radial length (PAR) represents priority adherence, while the angular width ($1-|\text{CB}|$) indicates bias resistance. Larger sectors indicate better priority control with less bias.
Most models fail to enforce instruction hierarchies consistently, as reflected in their small total areas. GPT-4o and GPT-4o-mini perform best, particularly in categorical constraints (language, case), likely due to their explicit instruction hierarchy training. However, even these models show significant variation across constraints, suggesting that their prioritization ability remains inconsistent.
@@
In general, models have better priority control over categorical constraints (e.g., case, language) than constraints requiring reasoning along a continuous scale (e.g., keeping counts during generation). This suggests that the limited priority control fails to generalize to more complex constraints.
@@
These findings reinforce that LLMs lack a robust mechanism for enforcing instruction priorities across diverse constraints, and also highlight a fundamental limitation in current instruction tuning paradigms.
Model-specific Constraint Biases
@@
Figure 4: Constraint Bias (CB) across six constraint dimensions. Positive values favor the right-side constraint, while negative values favor the left-side constraint, with magnitude reflecting bias strength. The highlighted zone shows shared biases.
@@
Constraint Bias (CB) scores reveal that models exhibit strong inherent preferences when resolving conflicting instructions, often overriding designated priority structures. Figure 4 visualizes these biases, where each bar cluster represents a constraint pair, and bars indicate model-specific tendencies. Most models display strong but inconsistent biases across constraint types. Bias magnitudes often exceed 0.5, indicating a clear default tendency toward certain constraints.
@@
Notably, some biases are widely shared across models. All models favor lowercase over uppercase text, prefer generating texts with more sentences, and tend toward avoiding keywords. This consistency across different model architectures suggests these biases might stem from common patterns in pre-training data or fundamental architectural designs in current models. For instance, the preference for lowercase likely reflects the predominance of lowercase text in training corpora.
@@
Despite these shared biases, other preferences vary sharply across models. Word length preferences are particularly diverse: Qwen-7B strongly favors shorter texts (<<50 words), while Llama-8B heavily prefers longer texts (>>300 words). Language choice and keyword usage frequency similarly show model-specific variations, suggesting these aspects are likely more influenced by individual architectural choices and training approaches than by mutual patterns in the data.
Latent Hierarchical Priors
@@
Our findings in previous sections show that system/user separation fails to enforce consistent instruction-following behavior when constraint conflicts are present. While models can resolve simple constraints independently, the presence of competing instructions reveals a failure to generalize the intended priority of system over user inputs. This breakdown suggests that the system and user roles, introduced primarily during post-training via instruction tuning or safety alignment, do not form a robust internalized hierarchy.
@@
Unlike the artificial system/user roles introduced during post-training, many forms of social and institutional hierarchy are deeply embedded in the natural language corpora used for pre-training. These relational patterns occur frequently and consistently across diverse domains, potentially enabling models to internalize them as latent inductive biases.
@@
Thus, the failure raises an important question: while models struggle to enforce priority based on the explicit system/user separation, might they instead exhibit greater sensitivity to “societal” hierarchies that are implicitly learned during pretraining? Recent jailbreak techniques exploit such social framings to override safety mechanisms (zeng-etal-2024-johnny). They operate on the implicit assumption that LLMs acquire social patterns during pretraining, which in turn may show stronger influence over model behavior than explicitly imposed guardrails. In this section, we examine whether such societal hierarchies function as stronger determinants of priority than system/user role designation.
@@
We examine three representative types of societal hierarchies in their simplest forms:
@@
All constraints are embedded within a single user message. Authority is indicated solely through minimal social framing (e.g., “CEO requires…” vs. “Intern requires…”). We use the same underlying tasks, constraints, and evaluation metrics described in previous sections: models are prompted using the same base tasks followed by the same pairs of conflicting constraints. We evaluate three models spanning a range of system/user prioritization performance (Priority Adherence Rate, PAR): Qwen (14.4%), Claude (23.6%), and GPT4o-mini (47.5%). Each hierarchy is tested under four configurations that vary both the ordering and the authority assignment of the two constraints. This design isolates the effects of authority framing from simple positional biases. [^3] Ordering is shown to have minimal influence in the Technical Appendix.
@@
Table 4: Priority Adherence Rate (PAR) to dominant constraint across societal hierarchy types vs. system/user separation.
Societal Hierarchies Induce Stronger Control Biases
Table 4 shows that societal hierarchy framings consistently yield higher priority adherence than the explicit system/user separation across models. For GPT4o-mini, PAR increases from 47.5% under system/user prompting to 77.8% when the dominant constraint is framed through social consensus, and Qwen-7B even rises from 14.4% to 65.8% under the same comparison. These patterns indicate that model priorities are more strongly shaped by familiar societal hierarchies than by artificial system/user markers.